# Web Privacy based on Contextual Integrity: Measuring the Collapse of Online Contexts

### v1

@@ 3.2. Case Selection
We test persistent browser identification trends across the top 100 popular websites in seven distinct online contexts (700 websites overall) - News & Media, Health, Education, Finance, LGBTQ, eCommerce, and Adult. Each context is constructed based on different purposes and values, carry different weights and identities for users and can be uniquely tagged to a distinct context: User interactions with news & media websites are based on the willingness to consume updated information or knowledge. Healthcare websites provide users with medical advice and serve as a space for users to advance their health. Educational websites signal users’ interest and educational aspirations. Users’ interactions with financial websites facilitate knowledge on efficient allocation of capital. Gender-related websites promote users’ gender interests and knowledge. eCommerce websites enable users to acquire products and services they would like to consume, and adult websites signal users’ sexual desires. These distinct contexts should not be exposed to trackers (=collapse within context) and should not be linked to other contexts from the same user (=collapse between contexts).
Following our selection of categories, we selected the top 100 popular websites from each of the seven categories (700 websites overall) based on the Similarweb.com list of popular websites. This provided a comprehensive cross-section list of widely visited sites across different contexts. Table 1 below indicates the average number of visitors from the top 100 websites in each context category. The Appendix section provides the full list of websites crawled for each context. We chose websites based on their popularity to learn about context collapse in the commercial Web, where we already know that other tracking patterns pervasively take place.
@@ 4. Data collection
Through an open-source instrumented Firefox browser implementation via OpenWPM, that simulates browser activity and records website responses, metadata, cookies used, and scripts executed (Englehardt and Narayanan, 2016), we studied the top 700 popular websites across seven online contexts that were crawled on a daily basis. We ran seven crawls per day for the 700 websites, starting with a different context every time to learn how user identification is further diffused by trackers from the first context to other contexts. Each context was a priority context and crawled first, followed by the other categories in a random order. We ran seven crawls of the selected 700 websites per day, for 28 days, From Oct 26 - November 28, from an AWS server located at Ashburn, Virginia, USA.
We studied stateful and stateless tracking practices via HTTP & JavaScript cookies and JavaScript APIs. For each experiment, we are matching cookie IDs among contexts to realize which trackers and from which websites the user IDs is diffused across contexts. As opposed to previous studies, we do not assume that the presence of trackers in two different contexts means that they persistently identify users across those contexts (Hu and Sastry, 2020b). Instead, we looked for a valid evidence, i.e. the usage of the same cookie-id across contexts, to assume persistent user identification. This leaves us only with trackers who engaged in uniquely identifying users’ browsers.
To improve the accuracy of our data collection efforts and ensure that the crawled websites are revealing their ’true’ behavior rather than minimizing their tracking practices due to recognition of our bot crawler, we implemented several recommended modifications to OpenWPM from Krumnow et al. (2022) (Krumnow et al., 2022):
  • Stealth JavaScript Instrumentation: By default, many websites can detect OpenWPM and similar tools, which might result in them adjusting their behavior. To counter this, we enabled a stealth mode for the JavaScript instrumentation, allowing our automated browser to be less detectable.
@@
  • Native Display Mode: Instead of using the more common ”headless” browser mode (where the browser runs without a graphical interface), we opted for a ”native” display mode with a full graphical user interface. This approach makes the browsing session appear more like a real user and decreases the likelihood of detection.
  • Dynamic Window Resolutions: We randomized window and screen resolutions across different browsing sessions, emulating the diverse range of devices used by actual users. This technique added another layer of trust to the browsing behavior of our bot, making it harder for websites to detect automation.
@@ 4.1. Collecting Patterns of Persistent Browser Identification and Context Collapse
During the seven crawling experiments per day, we collected both stateful and stateless browser identification patterns. For stateful browser identification via cookies, we collected two types of cookies: (1) HTTP Cookies - these are traditional cookies set by servers through HTTP headers using the Set-Cookie directive. These cookies are typically stored in the browser and sent with every subsequent HTTP request to the same domain; (2) JavaScript Cookies - these cookies are set or manipulated directly by a JavaScript on the client side, typically via the document.cookie API. Unlike HTTP cookies, JS cookies can be fully controlled by the website’s JavaScript, enabling dynamic interactions with the user session. Based on previous studies, we identified an ID cookie by the following four elements (Yang and Yue, 2020; Acar et al., 2014; Englehardt and Narayanan, 2016):
@@
  • Parameter-Value remains similar within an crawl.
@@
Both stateful and stateless browser identifiers were collected from the first crawled context, and then examined across subsequent crawls in other contexts. The crawl of 700 websites, seven times a day, was conducted over the period of 28 days to test the validity of our findings and tests whether differences between contexts are statistically significant. The data collection and analysis processes are visualized in figure 2.
@@ 5. Limitations
  • Crawling only homepages: We only crawled the homepages of selected websites to avoid bot detection. It is known that most tracking is happening on inner rather than homepages of websites (Englehardt and Narayanan, 2016; Urban et al., 2020), but at the same time, previous work found that crawling inner pages increases the chances of OpenWPM discovery by the crawled website (Krumnow et al., 2022).
@@
  • Not considering cookie-syncing practices: In our study of stateful user identifiers, we have not examined how trackers share user IDs via cookie syncing practices. User ID sharing between trackers is another information flow that leads to context collapse. Even if different user IDs are assigned for different websites, in different contexts, the linkage of IDs can be observed via HTTP requests in the browsing session (Papadopoulos et al., 2019). We did not include such inspection in our data analysis at the moment and plan to do so in a follow-up study.
  • Tracking pixels currently not under study: We have not studied the usage of tracking pixels across websites by trackers. Tracking pixels were found to be responsible for cookie syncing or for the purpose of learning about an already existing user identifier (Fouad et al., 2020). As part of our efforts to include cookie syncing practices to learn about context collapse, we plan to study cookie syncing through invisible pixels.
@@
  • Crawling top-popular websites: We chose to crawl top-popular websites, with tens of millions of user visits per website per month, to learn how context collapse happens in the popular Web. At the same time, we are not measuring context collapse based on users’ real browsing behavior. This requires a dedicated browser extension and agreement from users to share their data. We aim to develop this in future research efforts.
@@ 6.1. Multiple context collapse: Persistent user identification between contexts
Table 2 presents average statistics of multiple context collapse across different context origin. The first row, for example, describes the crawl of top 700 popular websites when the top-100 popular adult websites were crawled first. On average, 285.76 unique third-parties appeared in the adult context, and an average number of 461.8 cookies were dropped on the browser. 56.8 third-parties (19.88 percent of all third-parties in the adult context) were labeled as ’persistent identifiers’ because they persistently identified the browser from the adult context to one or more of the following contexts in that crawl. 9.04 percent of all crawled adult websites, on average, enabled persistent browser identification from their websites to other contexts. Differences between the mean of persistent identifiers from each context and participating websites from each context were found to be statistically significant based on single-factor Anova tests as detailed in the appendix.
@@
Figures 3 & 4, from different crawls that started from a different context illustrate this difference. In the visualized crawl in figure 3, from Oct 26, 2024, 60.49 percent of persistent identifiers originated from the adult context traveled all the way to the last context - LGBTQ. 18.52 percent of the identifiers stopped after six contexts (including the original one). The rest of the identifiers (less than 20 percent in total) diffused to three or less subsequent contexts. For persistent identifiers originated in LGBTQ websites, however, illustrated in figure 4, the diffusion is more even, making the figure more colorful. 20.78 percent traveled all the way to the finance context. The majority of identifiers (27.7 percent) traveled through six contexts (including the origin), 21.05 percent of the identifiers traveled through five contexts (including the origin). The rest 30 percent of identifiers stopped after 3 contexts or less.
@@
We now move to inspect the coverage of websites by the trackers that act as persistent identifiers. Previous work found a long-tail distribution for third-parties across the Web, with the vast majority of third-parties cover less than one percent of the websites that were studied (Sørensen and Kosta, 2019; Englehardt and Narayanan, 2016). In contrast, we found that persistent identifiers coverage is a short tail distribution across the websites they connect between contexts. Figure 5 shows the different distribution for each context, based on different context of origin. From six contexts, the highest percentage of websites covered by one tracker is 20 percent, while the rest of the trackers cover 0-10 percent of the websites across contexts. The adult context is an outlier, with a few trackers cover 30-45 percent of the connected websites from this context. This finding shows that persistent identification across multiple contexts does not follow known tracking trends that were found to be more centralized and dominated by a few trackers.
@@
For each context, there are a few dominant trackers and notable connected websites worth highlighting. Figure 6 visualizes multiple context collapse from the finance context. Trackers that are labeled as persistent identifiers are in the middle, connecting colored nodes that represent websites. The size of the nodes change based on their in- and out-degree. The crawling took place clock-wise, with the 12 o’clock context crawled first. We can see how persistent identification takes place by many actors (187) and connecting more websites between contexts (120), relative to those numbers in other contexts. Specifically, (foxbusiness.com), (economictimes.com), (xe.com), and (toyokeizai.net) are the financial websites that mostly connect browser identification to other contexts. They are notably connected to (ratemyprofessors.com) in the education context, (ssg.com) and (slickdeals.net) from the eCommerce context, (businessinsider.com) and (oneindia.com) in the News & Media context, (autostraddle.com) and (queerty.com) in the LGBTQ context, and (chemistwarehouse.com.au) in the health context. These sites are key participants in the collapse of online contexts for users originating from the finance context. The dominant trackers, in the middle of the figure, that are linking websites from the finance context to other contexts are (bing.com) that links 13 websites, (taboola.com) that links 13 websites, (demdex.net) that links 12 websites, and (rubiconproject.com) that links 12 websites. As shown in the distribution of persistent identifiers across websites (figure 5), the 187 persistent identifiers from the finance context almost evenly distributed across the websites they connect.
@@
Surprisingly, a few spotted persistent identifiers serve as first-party websites in our top-700 popular websites list, significantly increasing their visibility on users. The list of those websites, their first-party web context, and their third-party tracking contexts is in the appendix.
6.2. Single Context Collapse: Persistent user identification within-contexts
Table 4 presents average statistics of within-context collapse by trackers per context over 25 days of collected data. The second row, for example, describes average results from all crawls of top 700 popular websites when the top-100 popular eCommerce websites were crawled first. 597 unique third-parties appeared in the eCommerce context on average, and an average number of 1276 cookies were dropped on the browser. 102 third-parties (17.1 percent of all third-parties in the eCommerce context) were labeled as ’persistent identifiers’ because they persistently identified the browser across sites within the eCommerce context. 58.4 percent of all crawled eCommerce websites, on average, enabled persistent browser identification to other websites within that context.
@@ 6.2. Single Context Collapse: Persistent user identification within-contexts
Figure 7 shows the distribution of persistent identifiers within contexts across the websites they connect. Results are different from what was found in the analysis of multiple context collapse. Within contexts, we found a long tail distribution of trackers - a few trackers connect most of the websites within the same context. The adult context, which had a longer tail distribution in the between-contexts analysis, is now an outlier, having a shorter tail distribution. The leading persistent identifiers within contexts are the ’usual suspects’ from the advertising industry - (doubleclick.net) is the leading tracker within all contexts, with more or close to 50 percent visibility of websites within the contexts. (adnxs.com), and (criteo.com) are also popular within the News & Media and LGBTQ contexts. In the health context we see (adsrvr.org) as another leading context collapse initiator. In finance it is also (bing.com), in education we also see (facebook.com) and (linkedin.com), and in adult websites (yandex.ru) and (tsyndicate.com) persistently identify browsers across 10 and 7 percent of the websites.
@@
Figure 8 shows persistent browser identification within the education context. 71 percent of websites within this context have an out-degree in the network that is greater than one - meaning that the browser is persistently identified between those websites and other websites in that context. The leading websites with the highest out-degree are (asu.edu - 66), (colorado.edu - 64), and (tophat.com - 61). These websites are linking more than 60 percent of the websites crawled in this context. On the other end, the three websites with the highest in-degree are (gmu.edu - 64), (sydney.edu.au - 63), and (purdue.edu - 62). These are websites that are get connected from more than 60 percent of the websites in the education context.
@@
The trackers that frequently connect websites within the education context are the leading, well-known trackers from the advertising industry. Figure 9 visualizes their work, as the size of the nodes is determined by their degree. The four largest nodes are (doubleclick.net - connecting 57 sites), (facebook.com - connecting 34 sites), (linkedin.com - connecting 32 sites), and (youtube.com - connecting 27 sites).

### v2

@@ 3.2. Case Selection
We test persistent browser identification trends across the top 100 popular websites in seven distinct online contexts (700 websites overall) - News & Media, Health, Education, Finance, LGBTQ, eCommerce, and Adult. Each website category is constructed based on different purposes and values, carry different weights and identities for users, and can be uniquely tagged to a distinct context: User interactions with news & media websites are based on the willingness to consume updated information or knowledge. Healthcare websites provide users with medical advice and serve as a space for users to advance their health. Educational websites signal users’ interest and educational aspirations. Users’ interactions with financial websites facilitate knowledge on efficient allocation of capital. Gender-related websites promote users’ gender interests and knowledge. eCommerce websites enable users to acquire products and services they would like to consume, and adult websites signal users’ sexual desires. These distinct contexts should not be exposed to trackers (=collapse within context) and should not be linked to other contexts from the same user (=collapse between contexts).
Following our selection of categories, we selected the top 100 popular websites from each of the seven categories (700 websites overall) based on the similarweb.com list of popular websites. This provided a comprehensive cross-section list of heavily visited sites across different contexts. Table 1 below indicates the average number of visitors from the top 100 websites in each context category. The Appendix section provides the full list of websites crawled for each context. We chose websites based on their popularity to learn about context collapse in the commercial Web, where we already know that tracking patterns pervasively take place.
@@ 4. Data collection
Through an open-sourced instrumented Firefox browser, OpenWPM, that simulates browser activity and records website responses, metadata, cookies used, and scripts executed (Englehardt and Narayanan, 2016), we studied the top 700 popular websites across seven Web contexts. We ran seven crawls per day for the 700 websites, starting with a different context every time, to learn how user identification is further diffused by trackers from the first context to other contexts. Each context was a priority context and crawled first, followed by the other categories in a random order. We ran seven crawls of the selected 700 websites per day, for 28 days, From Oct 26 - November 28 of 2024, from an AWS server located at Ashburn, Virginia, USA.
We studied stateful and stateless tracking practices via HTTP & JavaScript cookies and JavaScript APIs. For each experiment, we are matching cookie IDs among contexts to realize which trackers, and from which websites, user IDs are diffused across contexts. As opposed to previous studies, we do not assume that the presence of trackers in two different contexts means that they persistently identify users across those contexts (Hu and Sastry, 2020b). Instead, we looked for a valid evidence, i.e. the usage of the same cookie-id across contexts, to assume persistent user identification. This leaves us only with trackers who engaged in uniquely identifying users’ browsers.
To improve the accuracy of our data collection efforts and ensure that the crawled websites are revealing their ’true’ tracking behavior to our bot crawler, we implemented several recommended modifications to OpenWPM from Krumnow et al. (2022) (Krumnow et al., 2022):
  • Stealth JavaScript Instrumentation: By default, many websites can detect OpenWPM and similar tools, which might result in them adjusting their behavior. In response, we enabled a stealth mode for the JavaScript instrumentation, allowing our automated browser to be less detectable.
@@
  • Native Display Mode: Instead of using the more common ”headless” browser mode, we opted for a ”native” display mode with a full graphical user interface. This approach makes the browsing session appear more like a real user and decreases the likelihood of detection.
  • Dynamic Window Resolutions: We randomized window and screen resolutions across different browsing sessions, emulating the diverse range of devices used by actual users. This technique added another layer of trust to the browsing behavior of our bot.
@@ 4.1. Collecting Patterns of Persistent Browser Identification and Context Collapse
During the seven crawling experiments per day, we collected both stateful and stateless browser identification patterns. For stateful browser identification via third-party cookies, we collected two types of cookies: (1) HTTP Cookies - traditional cookies set by servers through HTTP headers using the Set-Cookie directive. These cookies are typically stored in the browser and sent with every subsequent HTTP request to the same domain; (2) JavaScript Cookies - these cookies are set or manipulated directly by a JavaScript on the client side, typically via the document.cookie API. Unlike HTTP cookies, JS cookies can be fully controlled by the website’s JavaScript, enabling dynamic interactions with the user session. Based on previous studies, we identified an ID cookie by the following four elements (Yang and Yue, 2020; Acar et al., 2014; Englehardt and Narayanan, 2016):
@@
  • Parameter-Value remains similar within a crawl.
@@
Both stateful and stateless browser identifiers were collected from the first crawled context, and then examined across subsequent crawls in other contexts. The crawl of 700 websites, seven times a day, was conducted over the period of 28 days to test the validity of our findings and check whether differences in the quantity of tracking between contexts are statistically significant. The data collection and analysis processes are visualized in Figure 2.
@@ 5. Limitations
  • Crawling only homepages: We only crawled the homepages of selected websites to avoid bot detection. It is known that most tracking is happening on inner rather than homepages of websites (Englehardt and Narayanan, 2016; Urban et al., 2020), but at the same time, previous work found that crawling inner pages increases the chances of OpenWPM discovery (Krumnow et al., 2022).
@@
  • Not fully considering cookie-syncing practices: In our study of stateful user identifiers, we conducted a preliminary analysis of how trackers share user IDs via cookie syncing and forwarding practices, including through the use of tracking pixels. ID sharing between trackers is another information flow that leads to context collapse. Even if different user IDs are assigned for different websites, in different contexts, the linkage of IDs can be observed via HTTP requests in the browsing session (Fouad et al., 2020; Papadopoulos et al., 2019). Our preliminary analysis reveals that sharing of user IDs varies significantly between contexts, across both the number of trackers and participating websites. We spot interesting relations between trackers who already identified the user’s browser and trackers that did not directly drop an ID cookie on our browser. We plan to conduct an in-depth follow-up study in a separate paper about the impact cookie syncing and forwarding have on context collapse.
@@
  • Crawling top-popular websites: We chose to crawl top-popular websites, with tens of millions of user visits per website every month (see Table 1), to learn how context collapse happens in the popular Web. At the same time, we are not measuring context collapse based on users’ real browsing behavior. This requires a dedicated browser extension and agreement from users to share their data. We aim to develop this in future research efforts.
@@ 6.1. Multiple context collapse: Persistent user identification between contexts
Table 2 presents average statistics of multiple context collapse across different context origin. The first row, for example, describes the crawl of top 700 popular websites when the top-100 popular adult websites were crawled first. The first row indicates that on average, 285.76 unique third-parties appeared in the adult context, and an average number of 461.8 cookies were dropped on the browser. 56.8 third-parties (19.88 percent of all third-parties in the adult context) were labeled as ’persistent identifiers’ because they persistently identified the browser from the Adult context to one or more of the following contexts in that crawl. 9.04 percent of all crawled adult websites, on average, enabled persistent browser identification from their websites to other contexts. Differences between the mean of persistent identifiers from each context and participating websites from each context were found to be statistically significant based on single-factor ANOVA tests as detailed in the appendix.
@@
Figures 3 & 4, from different crawls that started from a different context, illustrate this difference. In the visualized crawl in Figure 3, 60.49 percent of persistent identifiers that originated from the Adult context traveled all the way to the last context - LGBTQ. 18.52 percent of the identifiers stopped after six contexts (including the original one). The rest of the identifiers (less than 20 percent in total) diffused to three or less subsequent contexts. For persistent identifiers originated in LGBTQ websites, however, illustrated in Figure 4, the diffusion is more even, making the visualization more colorful. 20.78 percent traveled all the way to the websites in the Finance context. The majority of identifiers (27.7 percent) traveled through six contexts (including the origin), 21.05 percent of the identifiers traveled through five contexts (including the origin). The rest 30 percent of identifiers stopped after 3 contexts or less.
@@
We now move to inspect the coverage of websites by trackers that act as persistent identifiers. Previous work found a long-tail distribution for third-parties across the Web, with the vast majority of third-parties cover less than one percent of the websites that were studied (Sørensen and Kosta, 2019; Englehardt and Narayanan, 2016). In contrast, we found that persistent identifiers coverage holds a short tail distribution across the websites they connect between contexts. Figure 5 shows the different distribution for each context, based on different context of origin. From six contexts, the highest percentage of websites covered by one tracker is 20 percent, while the rest of the trackers cover 0-10 percent of the websites across contexts. The adult context is an outlier, with a few trackers cover 30-45 percent of the connected websites from this context. This finding shows that persistent identification across multiple contexts does not follow known tracking trends that were found to be more centralized and dominated by a few trackers.
@@
For each context, there are a few dominant trackers and notable connected websites worth highlighting. Figure 6 visualizes multiple context collapse from the Finance context. Trackers that are labeled as persistent identifiers are in the middle, connecting colored nodes that represent websites. The size of the nodes change based on their in- and out-degree. The crawling took place clock-wise, with the 12 o’clock context crawled first. We can see how persistent identification takes place by many actors (187) and connecting more websites between contexts (120), relative to those numbers in other contexts. Specifically, (foxbusiness.com), (economictimes.com), (xe.com), and (toyokeizai.net) are the financial websites that mostly connect browser identification to other contexts. They are notably connected to (ratemyprofessors.com) in the education context, (ssg.com) and (slickdeals.net) from the eCommerce context, (businessinsider.com) and (oneindia.com) in the News & Media context, (autostraddle.com) and (queerty.com) in the LGBTQ context, and (chemistwarehouse.com.au) in the health context. These sites are key participants in the collapse of online contexts for users originating from the Finance context. The dominant trackers, in the middle of the figure, that are linking websites from the finance context to other contexts are (bing.com) that links 13 websites, (taboola.com) that links 13 websites, (demdex.net) that links 12 websites, and (rubiconproject.com) that links 12 websites. As shown in the distribution of persistent identifiers across websites (Figure 5), the 187 persistent identifiers from the finance context are almost evenly distributed across the websites they connect.
@@
Surprisingly, a few spotted persistent identifiers serve as first-party websites in our top-700 popular websites list, significantly increasing their visibility on users. The list of those websites, their first-party web context, and their third-party tracking contexts are detailed in the appendix.
6.2. Single Context Collapse: Persistent user identification within the context
Table 4 presents average statistics of within-context collapse by trackers per context, over 25 days of collected data. The second row, for example, describes average results from all crawls of top 700 popular websites when the top-100 popular eCommerce websites were crawled first. 597 unique third-parties appeared in the eCommerce context on average, and an average number of 1276 cookies were dropped on the browser. 102 third-parties (17.1 percent of all third-parties in the eCommerce context) were labeled as ’persistent identifiers’ because they persistently identified the browser across sites within the eCommerce context. 58.4 percent of all crawled eCommerce websites, on average, enabled persistent browser identification to other websites within that context.
@@ 6.2. Single Context Collapse: Persistent user identification within the context
Figure 7 shows the distribution of persistent identifiers within contexts across the websites they connect. Results are different from what was found in the analysis of multiple context collapse. Within contexts, we found a long tail distribution of trackers - a few trackers connect most of the websites within the same context. The adult context, which had a longer tail distribution in the between-contexts analysis, is an outlier again, having a shorter tail distribution. The leading persistent identifiers within contexts are the ’usual suspects’ from the advertising industry - (doubleclick.net) is the leading tracker within all contexts, with more or close to 50 percent visibility of websites within contexts. (adnxs.com), and (criteo.com) are also popular within the News & Media and LGBTQ contexts. In the Health context we see (adsrvr.org) as another leading context collapse initiator. In Finance it is also (bing.com), in Education we also see (facebook.com) and (linkedin.com), and in Adult websites (yandex.ru) and (tsyndicate.com) persistently identify browsers across 10 and 7 percent of the websites.
@@
Figure 8 shows persistent browser identification within the education context. 71 percent of websites within this context have an out-degree in the network that is greater than one - meaning that the browser is persistently identified between those websites and other websites in that context. The leading websites with the highest out-degree are (asu.edu - 66), (colorado.edu - 64), and (tophat.com - 61). These websites are linking more than 60 percent of the websites crawled in this context. On the other end, the three websites with the highest in-degree are (gmu.edu - 64), (sydney.edu.au - 63), and (purdue.edu - 62). These are websites that get connected from more than 60 percent of the websites in the education context.
@@
The trackers that frequently connect websites within the education context are the leading, well-known trackers from the advertising industry. Figure 9 visualizes their work. The size of the nodes is determined by their degree. The four largest nodes are (doubleclick.net - connecting 57 sites), (facebook.com - connecting 34 sites), (linkedin.com - connecting 32 sites), and (youtube.com - connecting 27 sites).